OpenStudy (anonymous):

Write the equation of a parabola with a vertex at (5, 2) and a directrix y = 1??

OpenStudy (anonymous):

first do you know what this would look like?

OpenStudy (anonymous):

because the directrix is a horizontal line and the vertex lies above the directrix, it will look something like this 

OpenStudy (anonymous):

that means the \(x\) part is squared and so you will have something that looks like \[4p(y-k)=(x-h)^2\] where the vertex is \((h,k)\) so in your case it is going to be \[4p(y-2)=(x-5)^2\]

OpenStudy (anonymous):

now all you need is \(p\) which is the distance between the vertex and the directrix

OpenStudy (anonymous):

From my notes, this is how the equation is supposed to be set up

OpenStudy (anonymous):

the distance between \((5,2)\) and \(y=1\) is 1, so \(p=1\) and the equation should be \[4(y-2)=(x-5)^2\]

OpenStudy (anonymous):

no matter, divide both sides by 4 to get what you want

OpenStudy (anonymous):

ok, that makes more sense now, thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

lets check the answer \(y-2=\frac{1}{4}(x-5)^2\) and see if it is right http://www.wolframalpha.com/input/?i=parabola+%28y-2%29%3D1%2F4%28x-5%29^2 yup looks good